OpenStudy (anonymous):

Simplify using the distributive property: 10x^3 −8x^2 −12x divde by 2x A. 5x 2 −4x−6 B. 5x 6 −4x 4 −6x 4 C. 5x 3 −4x 3 −6x D. 5x 4 −4x 2 −6

OpenStudy (anonymous):

Please check the question it last term @Jay-Bird00

OpenStudy (anonymous):

@dpasingh Huh?

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.

OpenStudy (anonymous):

\[\frac{10x^3 −8x^2 −12x}{ 2x} =\frac{10x^3 }{ 2x}−\frac{8x^2}{ 2x} −\frac{12x}{ 2x}\] \[=\frac{10 }{ 2}x^{3-1}−\frac{8}{ 2}x^{2-1} −\frac{12}{ 2}x^{1-1}\] \[=5x^{2}−4x^{1} −6x^{0}=5x^{2}−4x −6 \times 1\] \[=5x^{2}−4x −6\] @Jay-Bird00

OpenStudy (anonymous):

@dpasingh Thank you

OpenStudy (anonymous):

\[\frac{ 10x^3 −8x^2 −12x }{ 2x }\] 10x^3 ------ = 5x^2 Divide coefficient, subtract exponents. 2x -8x^2 ------ = -4x 2x -12x ----- = -6 2x answer is : 5x^2 -4x - 6